OpenStudy (anonymous):

Write an equation of a line that passes through the given point and is perpendicular to the given line in slope intercept form (4,-5) y=2x+3

jimthompson5910 (jim_thompson5910):

what is the slope of y = 2x+3

OpenStudy (anonymous):

2x

jimthompson5910 (jim_thompson5910):

close, it's just 2

jimthompson5910 (jim_thompson5910):

that's the same as 2/1

OpenStudy (anonymous):

my bad sorry.wouldnt it be -2/1 ??

jimthompson5910 (jim_thompson5910):

to find the perpendicular slope, just flip the fraction and flip the sign so flip the fraction to go from 2/1 to 1/2 then flip the sign to go from +1/2 to -1/2 this means that the perpendicular slope is -1/2

jimthompson5910 (jim_thompson5910):

now you need to find the equation of the line that a) has the slope of -1/2 and b) goes through the point (4, -5)

OpenStudy (anonymous):

AHHH!! I get it !!the new equation will be y=-1/2+3 now right?

jimthompson5910 (jim_thompson5910):

no, that's not correct

OpenStudy (anonymous):

no?

jimthompson5910 (jim_thompson5910):

you start with y = mx+b and you plug in the slope (-1/2) and the given point (4,-5) y = mx+b y = (-1/2)x + b ... plug in the slope m = -1/2 -5 = (-1/2)(4) + b ... plug in x = 4 and y = -5 -5 = -2 + b solve for b

OpenStudy (anonymous):

b=-3

jimthompson5910 (jim_thompson5910):

good, so the equation of the perpendicular line is \[\Large y = -\frac{1}{2}x - 3\]

OpenStudy (anonymous):

Thank you so much!:D

jimthompson5910 (jim_thompson5910):

you're welcome